What should you pay for a share of stock if the most recent dividend paid was $3.50 per share, you require a rate of return of 10%, and dividends are expected to grow at a constant rate of 2% per year?
Answer #1

Stock price = D1 / (interest rate - growth rate)

= 3.5 * 1.02 / (0.1 - 0.02)

= 3.5 * 1.02 / 0.08

= 44.625 ~ 44.63